One end of an insulated metal rod is maintained at 100 ∘C, while the other end is maintained at 0∘C by an ice-water mixture. The rod is 60.4 cm long and has a cross-sectional area of 1.25 cm2. The heat conducted by the rod melts 8.50 g of ice in 10.5 min.

Find the thermal conductivity k of the metal.
Answer #1

\frac{Q}{t}=\frac{KA\Delta T}{d}

Q=mL

L = latent heat of melting = 334 J/g

K=\frac{Qd}{tA\Delta T}

K=\frac{334*8.5*60.4*10^{-2}}{10.5*60*1.25*10^{-4}*100}=217.747\ Wm^{-1}K^{-1}